iA stock price is currently selling at sh. 50. It is known that at the end of six months it will be either sh. 45 or sh. 55. The risk-free rate is 10% per annum with continuous compounding. What is the value of a six-month European call option with a strike price of sh. 50?
1
Expert's answer
2016-08-15T08:42:03-0400
If a stock price is currently selling at sh. 50, at the end of six months it will be either sh. 45 or sh. 55 and the risk-free rate is 10% per annum with continuous compounding, then we have u = 1.1, d = 0.9, r = 0.10, T = 0.5, and K = 50.
So, p = (e^(0.10)(0.5) − 0.9)/(1.1 - 0.9) = 0.7564 and 1 − p = 0.2436.
The value of the put is therefore f = e^-(0.10)(0.5)*[0.7564*0 + 0.2436*5] = $1.16